OpenStudy (anonymous):

Solve 3 log6 2 + 2 log6 3 − log6 x = 2.

OpenStudy (anonymous):

a.x = 2 b.x = 6 c.x = 15 d.x = 36

OpenStudy (mathstudent55):

This is where all the log rules come into play.

OpenStudy (anonymous):

yes, logarithms give me a headache

OpenStudy (solomonzelman):

some rules of logarithms: \(\color{#000000 }{ \displaystyle \log_a(b)+\log_a(c)=\log_a(b\cdot c) }\) \(\color{#000000 }{ \displaystyle \log_a(b)-\log_a(c)=\log_a(b\div c) }\) \(\color{#000000 }{ \displaystyle w\log_a(b)=\log_a(b^w) }\)

OpenStudy (anonymous):

would it be x=6?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 3\log_6(2)+2\log_6(3)-\log_6x=2 }\) write it as a single logarithm and then use (another rule) \(\color{#000000 }{ \displaystyle \log_a(b)=c\quad \Longleftrightarrow \quad a^c=b }\)

OpenStudy (mathstudent55):

Use the rules of logs on the left side to end up with a single log of a number.

OpenStudy (solomonzelman):

You want an example, perhaps?

OpenStudy (anonymous):

yes please

OpenStudy (solomonzelman):

\(\color{#0000ff }{ \displaystyle 4\log_2(2)-2\log_2(x)+ 2\log_2(10)=2}\) after applying; \(w\log_a(b)=\log_a(b^w)\), we get: \(\color{#0000ff }{ \displaystyle \log_2(2^4)-\log_2(x^2)+ \log_2(10^2)=2}\) \(\color{#0000ff }{ \displaystyle \log_2(16)-\log_2(x^2)+ \log_2(100)=2}\) then, we'll use; \(\log_a(B)+\log_a(C)=\log_a(B\cdot C)\), and \(\log_a(B)+\log_a(C)=\log_a(B\cdot C)\) we get: \(\color{#0000ff }{ \displaystyle \log_2(16\div x^2\times 100)=2}\) or, \(\color{#0000ff }{ \displaystyle \log_2\left(\frac{1600}{x^2}\right)=2}\) And then we apply; \(\color{ }{ \displaystyle \log_a(b)=c\quad \Longleftrightarrow \quad a^c=b }\) \(\color{#0000ff }{ \displaystyle \left(\frac{1600}{x^2}\right)=2^{2}}\) and then it's just a algebra \(\color{#0000ff }{ \displaystyle \frac{1600}{x^2}=4}\) \(\color{#0000ff }{ \displaystyle \frac{1600}{4x^2}=1}\) \(\color{#0000ff }{ \displaystyle 1600=4x^2 }\) \(\color{#0000ff }{ \displaystyle 400=x^2\quad \Longrightarrow \quad x=\pm20 }\)

OpenStudy (anonymous):

I am so confused :(